OpenStudy (anonymous):

how do you solve: f(x)= x^2, [-2,1] with the mean value theorem

OpenStudy (anonymous):

i am going to make a guess that this means "show that f satisfies the hypothesis of the mean value theorem, and find the c for which \[f'(x)=\frac{f(1)-f(-2)}{3}\]

OpenStudy (anonymous):

That's a nice guess satellite. I'm sure you meant that \(f'(c)=\frac{f(1)-f(-2)}{3}\), where \(-2\le c \le 1\).

OpenStudy (anonymous):

yes that is what i meant. these questions always annoy me because it turns a theorem into a calculation, as if students are too dumb to understand and if - then statement. of course mvt doesn't say "solve for c"